# Golden Touchstone: A Comprehensive Bilingual Benchmark for Evaluating Financial Large Language Models

### v1

@@ 3.2 Golden Touchstone Benchmark Design
Task Dataset Train Valid Test Metrics
Sentiment Analysis FinFE-CN 16157 2020 2020 Weighted-F1 ACC
Classification FinNL-CN 7071 884 884 ORMACC
Entity Extraction FinESE-CN 14252 1781 1782 ORMACC
Relation Extraction FinRE-CN 13486 1489 3727 RMACC
Multiple Choice FinEval 1071 170 3340 Weighted-F1 ACC
CPA 6268 1444 6 Weighted-F1 ACC
Summarization FinNA-CN 28800 3600 3600 ROUGE BLEU
Question Answering FinQa-CN 19906 2469 2480 RMACC
FincQa-CN 21965 2741 2745 RMACC
Stock Movement Prediction AStock 11815 1477 1477 Weighted-F1 ACC
English Finance Benchmark: We constructed this benchmark primarily based on the PIXIU and FinBen Benchmark. For sentiment analysis, we selected the FPB (Malo et al., 2014) and FiQA-SA (Maia et al., 2018b) datasets, evaluated using accuracy (ACC) and weighted F1 score (weighted-F1), two widely adopted classification metrics. For classification tasks, we used the Headlines (Sinha and Khandait, 2021), FOMC (Shah et al., 2023), and LendingClub (Feng et al., 2023) datasets. Evaluation of the Headlines and FOMC datasets was conducted with ACC and weighted-F1, while the LendingClub dataset was assessed using Matthews Correlation Coefficient (MCC) and accuracy to align with FinBen standards. For entity extraction, we used the NER (Alvarado et al., 2015) dataset, employing entity-level F1 score (Entity-F1) in line with the PIXIU framework. The EDTSUM (Zhou et al., 2021) dataset was selected for summarization tasks, evaluated using Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) and ROUGE scores (Lin, 2004), which are standard metrics in text generation. For question answering, we chose the FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) datasets, opting for Regex Match Accuracy (RMACC) over exact match accuracy, as RMACC better accommodates our general-purpose models that may not directly output exact answers. For relation extraction, we used the FinRED (Sharma et al., 2022) dataset, evaluated with the Relation-F1 metric, consistent with FinGPT. The CFA (Yang et al., 2024b) dataset was used for multiple-choice tasks, evaluated using weighted-F1 and ACC. Finally, we selected the DJIA (Aaron7Sun, 2016) dataset for stock movement prediction, evaluated using ACC and weighted-F1. Unlike PIXIU’s use of MCC for this task, we employed weighted-F1 due to generation models’ tendency to produce low-accuracy [unknown] labels.
Chinese Finance Benchmark: We constructed this benchmark primarily based on the BBT-Benchmark (Lu et al., 2023a) and FinEval Benchmark (Zhang et al., 2023). For sentiment analysis, we used the FinFE-CN dataset, evaluated by weighted-F1 and ACC. For question answering, we used the FinQA-CN and FinCQA-CN datasets, assessed using RMACC. Classification tasks employed FinNL-CN, requiring full matches for multi-answer questions and evaluated by Overall Regex Match Accuracy (ORMACC).For relation extraction, we used the FinRE-CN dataset, also evaluated with RMACC. The entity extraction task employed the FinESE-CN dataset, evaluated using ORMACC due to the dataset’s lack of BIO labels for original text entities. Summarization tasks used the FinNA-CN dataset, evaluated with BLEU and ROUGE scores in line with the English benchmark. For multiple-choice tasks, we used FinEval and CPA (Yang et al., 2024b), with FinEval combining finance, economy, accounting, and certification questions, both evaluated with weighted-F1 and ACC. The AStock (Zou et al., 2022) dataset was used for stock movement prediction, assessed by ACC and weighted-F1.
In tasks with [unknown] labels, such as classification, multiple choice, and stock prediction, we treated [unknown] labels as incorrect to maintain consistency. This adjustment proved more reliable than alternative methods like sample omission, random assignment, or fixed selection, which may introduce bias.
4 Experiments
4.1 Experimental Setup
Baselines. We conducted an extensive experimental evaluation against the Golden Touchstone Benchmark, incorporating a comprehensive array of models. For all models and inference tasks, we set the PyTorch and CUDA random seeds and configured the model with a greedy decoding strategy. This ensures reproducibility of experimental results and eliminates the influence of sampling decoding strategies on the final generated outputs. This included cutting-edge commercial models such as GPT-4o (OpenAI, 2023), alongside prominent open-source alternatives like Meta Llama-3 (AI@Meta, 2024) and Alibaba Qwen-2 (Yang et al., 2024a). Additionally, we integrated the latest and most influential financial language models (FInLLMs), namely FinGPT (Yang et al., 2023), FinMA (Xie et al., 2024), CFGPT (Li et al., 2023), and DISC-FinLLM Chen et al. (2023). These models were meticulously selected to represent a diverse spectrum of capabilities, ranging from general-purpose language understanding to specialized financial domain expertise. Our experiments aimed to rigorously assess the performance, robustness, and adaptability of each model within the context of financial data processing and analysis. The results provide valuable insights into the strengths and limitations of current state-of-the-art models, offering a foundation for future advancements in financial language modeling.
Touchstone-GPT Training. To further contribute to the research and development of FInLLMs and Financial benchmarks for LLMs, we have meticulously trained and open-sourced a Touchstone-GPT model. This initiative aims to serve as a valuable resource for advancing the field, providing a robust and versatile model that can be utilized for a wide range of financial language tasks. By making the Touchstone-GPT model publicly available, we aspire to foster collaboration and innovation within the research community, facilitating the development of more sophisticated and effective financial language models. Our hope is that this open-source contribution will not only enhance the current state of research but also inspire new methodologies and benchmarks that push the boundaries of what is possible in financial language modeling. We adopted a two-stage training strategy comprising continuous pre-training and post-training, based on the Qwen-2 (Yang et al., 2024a) foundational model. During the continuous pre-training phase, we initially conducted pre-training on a high-quality financial corpus containing 100 billions tokens, which included textbooks, encyclopedias, research reports, news articles, and real-time analysis, all meticulously cleaned. In the post-training phase, we employed a standard instruction fine-tuning strategy, collecting, cleaning, and formatting a high-quality dataset of 300,000 instruction-response pairs shown in Tabel.3 and Table.4. To avoid catastrophic forgetting in general tasks, we also incorporated general-domain pre-training corpora (Gan et al., 2023) and instruction-tuning corpora (Peng et al., 2023) into continuous pre-training and post-training. This culminated in the final Touchstone-GPT model. We utilized Megatron(Shoeybi et al., 2019) for continuous pre-training and LlamaFactory(Zheng et al., 2024) for instruction post-training as our training frameworks, respectively. In this study, we employ an advanced model training setup using the AdamW optimizer (Kingma and Ba, 2014) with a learning rate of 1.0e-5, cosine annealing scheduler (Szegedy et al., 2016), and a 10% warmup ratio (He et al., 2016; Goyal et al., 2017) to enhance training stability and convergence. We enable gradient accumulation (Shoeybi et al., 2019) and checkpointing (Chen et al., 2016) to simulate larger batch sizes and reduce memory footprint. Training is conducted in mixed bfloat16 precision (Micikevicius et al., 2017; Wang and Kanwar, 2019) with DeepSpeed’s ZeRO-1 optimization (Rajbhandari et al., 2020), reducing memory consumption and allowing for larger model training. This comprehensive setup optimizes efficiency and performance, providing an effective solution for large-scale deep learning model training.
Our training was conducted on 4 NVIDIA DGX servers, each equipped with 8 A100 GPUs, and spanned a period of 4 weeks. Inference was performed on a single NVIDIA DGX server with eight A100 GPUs, utilizing parallel batch inference. During the pre-training phase, we employed a data packing strategy (Krell et al., 2021) and batch dynamic right padding strategy in the instruction tuning phase (Wolf et al., 2020), while the inference phase incorporated a batch left padding strategy (Wolf et al., 2020).
4.2 Evaluation Results
Figure 3: Comparison of different models’ performance across tasks in the Golden Touchstone benchmark, illustrating average performance for English and Chinese tasks respectively.
4.2.1 Results of English Benchmark
We analyze the English benchmark results shown in Table.5 across task categories shown in left column of Figure.3 such as Sentiment Analysis, Classification, and Summarization. We highlight specific strengths and weaknesses observed in each task and metric.
GPT-4o consistently performs well in tasks like Sentiment Analysis with a Weighted-F1 score of 0.8084 and an accuracy of 0.8093 on the FPB dataset and Multiple Choice, showcasing robustness in understanding sentiment-based and structured questions. However, it struggles significantly in Relation Extraction achieving a Relation-F1 score of 0.1613 on the FinRE dataset and Entity Extraction with an Entity-F1 score of 0.1800 on the NER dataset, suggesting difficulties with detailed information extraction and handling relationships. FinMA-7B excels in Sentiment Analysis on the FPB dataset, achieving a Weighted-F1 score of 0.9400 and an accuracy of 0.9402 but has difficulties in Classification on the FOMC dataset, with a Weighted-F1 score of 0.3988 and an accuracy of 0.4274 and Summarization on the EDTSUM dataset, with Rouge-1, Rouge-2, and BLEU scores all below 0.16, indicating a lack of versatility in broader NLP tasks despite its sentiment strengths. Qwen-2-7B-Instruct performs well in Sentiment Analysis with a Weighted-F1 score of 0.7965 on the FPB dataset but is weak in Summarization on the EDTSUM dataset, achieving a Rouge-2 score of 0.0433 and Question Answering with an RMACC of 0.0270 on the Finqa dataset. Llama-3-8B-Instruct stands out in Stock Movement Prediction achieving a Weighted-F1 score of 0.5116 but struggles in Entity Extraction achieving an Entity-F1 score of 0.2973, highlighting its capacity in structured tasks but limitations in text understanding. FinGPT-8B-lora shows consistent underperformance, especially in Relation Extraction (FinRE, Relation-F1: 0.0100) and Summarization (EDTSUM, BLEU: 0.0592), indicating significant gaps in training these complex financial tasks, necessitating targeted model improvements. The results of many models on lengdingclub are not good enough, this is because of the specialization of the dataset, if the model has not been trained on a similar dataset, the results from zero-shot inference are definitely not good, which emphasizes the importance of instruction fine-tuning for large models in finance. Touchstone-GPT, built on the strong foundation of Qwen-2-7B, utilizes high-quality financial pre-training and instruction tuning datasets through continuous pre-training and post-training, resulting in competitive performance across most tasks.
In summary, the analysis reveals that while GPT-4o and FinMA-7B perform well in sentiment tasks, Llama-3-8B-Instruct has strengths in stock movement prediction, but it’s still not at a practical level. All models, including Touchstone-GPT, have limitations in handling broader NLP tasks.
4.2.2 Results of Chinese Benchmark
We present an analysis of the Chinese benchmark results across tasks such as Sentiment Analysis, Entity Extraction, and Summarization. Specific task metrics are discussed to highlight strengths and weaknesses.
GPT-4o excels in Question Answering with an RMACC of 0.6578 on the FinQa-CN dataset and Entity Extraction achieving an ORMACC of 0.6867 on the FinESE-CN dataset but struggles in Classification with a score of 0.3303 on the FinNL-CN dataset and Relation Extraction with a score of 0.2754 on the FinRE-CN dataset, showing strength in detailed extraction but limitations in handling complex relationships. Qwen-2-7B-Instruct shows moderate but balanced performance, with strengths in Multiple Choice achieving a Weighted-F1 score of 0.7230 on the FinEval dataset but weaknesses in Relation Extraction achieving an ORMACC of 0.1330 and Entity Extraction achieving an ORMACC of 0.3678. Llama-3-8B-Instruct performs particularly well in Stock Movement Prediction (Weighted-F1: 0.4929) but is weak in Sentiment Analysis with a score of 0.4262 on the FinFe-CN dataset and Classification with a score of 0.0747, indicating a less capability in detailed financial text understanding. CFGPT1-7B-Full struggles in most tasks, particularly in Classification with a score of 0.0894 and Relation Extraction (0.0678), with low scores across Summarization metrics with a BLEU score of 0.0238 on the FinNA-CN dataset, implying a need for enhanced adaptability and specialized training. DISC-FinLLM-Full shows moderate potential in Entity Extraction with 0.4346 ORMACC but lacks in Relation Extraction (0.1182) and Classification (0.0011), suggesting opportunities for targeted optimization to enhance broader capabilities. As mentioned above, Touchstone-GPT, based on the Qwen-2-7B model, leverages carefully curated financial pre-training data and instruction tuning through continuous pre-training and post-training, achieving competitive results across most tasks.
In conclusion, while GPT-4o and Llama-3-8B-Instruct have areas of strength in Entity Extraction and Stock Movement Prediction, the models overall, including Qwen-2-7B-Instruct, CFGPT1-7B-Full, and DISC-FinLLM-Full, exhibit considerable gaps in handling relational and classification tasks in Chinese, highlighting the challenges of processing complex Chinese financial text. The need for targeted model improvements remains evident for these tasks.
4.3 Analysis of Evaluation Results
From the perspective of individual models, GPT-4o shows strong performance in sentiment analysis and structured tasks like multiple choice, indicating robust general language understanding capabilities. However, its weakness lies in relation extraction and detailed entity extraction, which require detailed understanding of complex financial relations. FinMA-7B stands out in sentiment tasks but lacks versatility, especially in question answering and summarization, likely due to the absence of targeted training for diverse NLP challenges. Qwen-2-7B-Instruct has a balanced yet modest performance, doing well in sentiment analysis but struggling significantly in question answering and summarization, which suggests a need for more specialized post-training. Llama-3-8B-Instruct excels in stock movement prediction, but shows limitations in tasks requiring deep linguistic understanding, such as entity and relation analysis. The metrics of FinGPT-8B-lora indicating that the current level of domain-specific tuning is insufficient for complex financial tasks. Finally, DISC-FinLLM-Full and CFGPT1-7B-Full demonstrate moderate strengths in entity extraction tasks but lack the robustness needed for broader NLP capabilities, revealing significant gaps in financial language comprehension.
From a task perspective, we observe that Sentiment Analysis generally yields high scores across most models, particularly for the English benchmark, indicating that sentiment understanding, even in financial contexts, is relatively well addressed by these models. In contrast, Relation Extraction and Question Answer in financial domain exhibit notably lower performance, especially for the Chinese benchmark. These results suggest that capturing financial relationships and classifying detailed financial statements pose greater challenges, requiring more sophisticated training datasets or better model architectures. The LendingClub dataset in Classification is a specialized dataset in the field of risk control, requiring more targeted fine-tuning to achieve good results. Stock Movement Prediction also shows low performance across most models, with only a few models such as Llama-3-8B-Instruct demonstrating relatively moderate performance, but it is still practically unusable, highlighting the inherent difficulty of this task. Market prediction relying solely on news information is likely insufficient; volume-price data and factor analysis can provide more comprehensive information. However, current large language models are unable to process these inputs, which is a significant area of future research. Summarization also stands out as a weak area for most models, with consistently low BLEU and Rouge scores, reflecting the challenges in generating concise, coherent summaries of complex financial text.
Table 5: Performance metrics of financial large language models across english tasks like Sentiment Analysis, Classification, and Summarization. Models include GPT-4o, Llama-3-8B, Qwen-2-7B, FinMA-7B, FinGPT-8B, and Touchstone-GPT. The best results of each dataset are marked in bold.
Task Dataset Metrics GPT-4o FinMA-7B full Qwen-2-7B Instruct Llama-3-8B Instruct FinGPT-8B lora Touchstone GPT
Sentiment Analysis FPB Weighted-F1 0.8084 0.9400 0.7965 0.7631 0.2727 0.8576
ACC 0.8093 0.9402 0.8000 0.7660 0.3072 0.8557
Fiqa-SA Weighted-F1 0.8106 0.8370 0.6726 0.7515 0.5885 0.8591
ACC 0.7702 0.8340 0.5957 0.7064 0.5872 0.8638
Classification Headlines Weighted-F1 0.7857 0.9739 0.7278 0.7006 0.4516 0.9866
ACC 0.7931 0.9739 0.7252 0.7004 0.4331 0.9866
FOMC Weighted-F1 0.6603 0.3988 0.6112 0.4904 0.2758 0.8788
ACC 0.6794 0.4274 0.6210 0.5625 0.2702 0.8790
lendingclub Weighted-F1 0.6730 0.1477 0.5938 0.5943 0.5480 0.9783
MCC 0.1642 -0.6218 0.1714 0.1670 -0.1120 0.9297
Entity Extraction NER Entity-F1 0.1800 0.6200 0.2875 0.2973 0.0231 0.6993
Relation Extraction FinRE Relation-F1 0.1613 0.0054 0.1083 0.0540 0.0100 0.5331
Multiple Choice CFA Weighted-F1 0.7700 0.2200 0.6697 0.5800 0.3993 0.7497
ACC 0.7700 0.2400 0.6700 0.5800 0.3800 0.7500
Summarization EDTSUM Rouge-1 0.1675 0.1566 0.1466 0.1467 0.0622 0.5254
Rouge-2 0.0556 0.0491 0.0433 0.0429 0.0085 0.3446
Rouge-L 0.1069 0.1060 0.0857 0.0930 0.0412 0.4705
BLEU 0.1192 0.1361 0.0999 0.1085 0.0592 0.4512
Question Answering Finqa RMACC 0.1037 0.0497 0.0270 0.0470 0.0110 0.2258
Convfinqa RMACC 0.2540 0.0953 0.0644 0.1477 0.0772 0.5053
Stock Movement Prediction DJIA Weighted-F1 0.4241 0.3211 0.2744 0.5116 0.2171 0.4396
ACC 0.4648 0.3291 0.4372 0.5101 0.2211 0.4749
Table 6: Performance metrics of financial large language models across chinese tasks like Sentiment Analysis, Classification, and Summarization. Models include GPT-4o, Llama-3-8B, Qwen-2-7B, CFGPT-7B, DISC-FinLLM, and Touchstone-GPT. The best results of each dataset are marked in bold.
Task Dataset Metrics GPT-4o Qwen-2-7B Instruct Llama-3-8B Instruct CFGPT1-7B Full DISC-FinLLM Full Touchstone GPT
Sentiment Analysis FinFe-CN Weighted-F1 0.6593 0.6274 0.3633 0.2528 0.4177 0.7888
ACC 0.6500 0.6436 0.4891 0.2732 0.4292 0.7936
Classification FinNL-CN ORMACC 0.3303 0.0622 0.0747 0.0894 0.0011 0.8360
Entity Extraction FinESE-CN ORMACC 0.6867 0.3678 0.3088 0.3863 0.4346 0.9074
Relation Extraction FinRE-CN RMACC 0.2754 0.1330 0.1296 0.0678 0.1182 0.6541
Multiple Choice FinEval Weighted-F1 0.7364 0.7230 0.4432 0.3543 0.4288 0.7361
ACC 0.7353 0.7235 0.4471 0.3529 0.4294 0.7353
CPA Weighted-F1 0.6312 0.6957 0.3421 0.3543 0.3451 0.9238
ACC 0.6309 0.6960 0.3504 0.3553 0.3518 0.9238
Summarization FinNA-CN Rouge-1 0.3197 0.3326 0.3477 0.1018 0.3486 0.5526
Rouge-2 0.1434 0.1597 0.1702 0.0263 0.1678 0.3603
Rouge-L 0.2511 0.2644 0.2802 0.0650 0.2997 0.5214
BLEU 0.1423 0.1541 0.1672 0.0238 0.1885 0.3944
Question Answering FinQa-CN RMACC 0.6578 0.5043 0.4540 0.1126 0.3949 0.9214
FinCQa-CN RMACC 0.4765 0.3422 0.3787 0.2714 0.2134 0.8552
Stock Movement Prediction AStock Weighted-F1 0.5007 0.4906 0.4903 0.4631 0.4142 0.4003
ACC 0.5017 0.4915 0.4956 0.4888 0.4144 0.5587
5 Conclusion and Future Work
Despite these advancements, the benchmark currently exhibits certain limitations, including a limited range of NLG tasks and a focus solely on single-modality. Future enhancements will include the integration of additional NLG tasks, such as extended text generation for financial report analysis and more sophisticated sentiment assessments. Furthermore, we plan to expand the benchmark to cover other financial sectors such as insurance, cryptocurrency, and futures trading, thus broadening the scope and applicability of financial LLM assessments across diverse scenarios. Also, the performance of Touchstone-GPT on specific tasks within the Golden Touchstone benchmark, particularly in stock market prediction, requires further improvement. Our subsequent research will explore the incorporation of agent-based and retrieval-augmented generation (RAG) methods to augment the model’s capabilities in numerical computation and real-time news analysis. Additionally, we aim to venture into multimodal modeling, integrating visual data and time-series data for tasks such as financial time-series forecasting, financial chart analysis, and content generation.
Appendix A Examples of instruction construction
This Table.7 provides a detailed overview of the instruction construction examples for various financial NLP tasks, categorized by task type and language. Each example includes the instruction, input, and output components, demonstrating the specific requirements and expected responses for each task. For a more comprehensive understanding of the prompt settings and additional details, please refer to our dataset’s open-source repository on Hugging Face: Golden-Touchstone Dataset.
Table 7: Examples of Instruction Construction for Various Financial Language Tasks, Categorized by Task Type and Language
Task Type Language Instruction Input Output
Sentiment Analysis English What is the sentiment of the following financial post: Positive, Negative, or Neutral? RT @tomhend777 \$MU needs to hold here -Broken for now. Needs big flush. Still not technically oversold so now big bounce yet neutral
Chinese 以下是股民论坛中的一则股民评论,其中包含有感性的情感输出和理性的涨跌预测等内容…… 判断的非常准确,几次T的相当稳妥! 1
Classification English Review the sentence from a central bank’s communiqué…… In their discussion of prices, participants indicated that data over the intermeeting period…… neutral
Chinese 把接下来输入的金融新闻分类为一个或多个与其描述内容相关的类别…… 加拿大皇家银行:将Affirm Holdings(AFRM.O)目标价从175美元下调至127美元。 外国 公司
Entity Recognition English In the sentences extracted from financial agreements in U.S. SEC filings…… There is a default in any agreement to which Borrower or any Guarantor is a party with a third party or parties…… Borrower, PER
Chinese 给定一段文本T,和文本所属的事件类型S,从文本T中抽取指定事件类型S的事件主体…… 文本: 天龙新材关联担保事项未及时披露被监管佳士科技(300193)股东减持900万股 套现近2亿 事件类型: 信批违规 天龙新材
Relation Extraction English What is the relationship between Ivan Glasenberg and Glencore in the context of the input sentence…… The persistent oversupply is "damaging the credibility of the industry," Glencore CEO Ivan Glasenberg said in May. owner_of
Chinese 给定句子和其中的头尾实体,要求你预测头尾实体之间的关系…… 头实体: ISIS 尾实体: 美军 句子: 美军已对ISIS发动<N>次空袭外资石油巨头欲撤离 unknown
Task Type Language Instruction Input Output
Stock Movement Prediction English Please predict the next rise or fall of DJIA Adj based on the next input of the day’s 25 most popular news items…… Top1:WikiLeaks demands answers after Google hands staff emails to US government…… 1
Chinese 在考量了公司的相关公告之后,请根据新闻对股票数据的影响对该公司股票的表现进行分类…… 公司董事长车成聚承诺自本公告日起未来六个月拟增持价值0.5-1.0亿公司股份…… 0
Multiple Choice English Given a text T, and several options, according to the question posed in the text T…… The inventory/sales ratio is most likely to be rising…… C
Chinese 给定一段文本T,和四个选项ABCD,根据文本T中提出的问题从四个选项中选择合适的多个选项作为答案…… 下列选项中责任中心判断一项成本是否可控的条件有( )…… A,B,D
Summarization English You are given a text that consists of multiple sentences…… PORTLAND, Ore., Feb. 17, 2021 /PRNewswire/ – Allied Market Research published a report, titled,"Matcha Tea Market By Product Type…… Matcha Tea Market to Reach \$4.48 Bn, Globally, by 2027 at 7.1%……
Chinese 请对根据接下来的输入的中文短新闻进行摘要总结,请直接开始总结,不需要输出任何解释 美港电讯APP 13日讯,法航荷航集团(Air France-KLM)已开始与波音(BA.N)和空客就可能成为该集团有史以来最大的飞机订单进行谈判…… 波音空客将竞争法航荷航集团史上最大订单
Question Answering English Please answer the given financial question based on the context…… on november 18 , 2014 , the company entered into a collateralized reinsurance agreement with kilimanjaro…… The answer is:0.26685
Chinese 请根据下面提出的一个问题,问题后的材料内会有相应的答案…… 江苏金沙地理信息股份有限公司公司上市事件对应的证券代码是什么?挖贝网10月9日,全国中小企业股转系统公告显示…… 873361
Appendix B Inference Template of Large Language Models
The Table.8 showcases how inference templates vary across different models. It is crucial to select the appropriate template for constructing correct inputs when inferring on the test sets of datasets. An incorrect template can significantly impair the performance of a model. We have observed that the underperformance of some large financial language models in some benchmarks is precisely due to not selecting the appropriate templates for evaluation. For more details of training and inference template, please refer to our open-source code repository on Github: Golden-Touchstone Code.

### v2

@@ 7 Ethical Statements
Task Dataset Train Valid Test Metrics License
Sentiment Analysis FinFE-CN 16157 2020 2020
Weighted-F1
ACC
Public
Classification FinNL-CN 7071 884 884
ORMACC
Public
Entity Extraction FinESE-CN 14252 1781 1782
ORMACC
Public
Relation Extraction FinRE-CN 13486 1489 3727
RMACC
Public
Multiple Choice FinEval 1071 170 3340
Weighted-F1
ACC
Apache-2.0 license
CPA 6268 1444 6
Weighted-F1
ACC
CC BY-SA 3.0
Summarization FinNA-CN 28800 3600 3600
ROUGE
BLEU
Public
Question Answering FinQa-CN 19906 2469 2480
RMACC
Public
FincQa-CN 21965 2741 2745
RMACC
Public
Stock Movement Prediction CMIN-CN 214873 23904 23571
Weighted-F1
ACC
MIT License
Appendix B Additional Experiment Results
Due to computational constraints imposed by our large-scale dataset (more than 112K samples; see Appendix A) and to ensure a fair comparison among models of similar scale, the main paper (Table 2 and Table 3) primarily reports results from models with <10B parameters, alongside GPT-4o. In this appendix, we supplement those findings with evaluations of several advanced or larger models (>100B parameters) in Table.6, providing a more comprehensive performance comparison. These models include Yi-34B(young2024yi), GLM-4-32B-0414(glm2024chatglm), Mixtral 8*7B(jiang2024mixtral), Qwen2.5-32B/72Byang2024qwen2, Xuanyuan 2.0(zhang2023xuanyuan) and deepseek-v3(liu2024deepseekv3). Furthermore, we extended the CMIN dataset to introduce Stock-CoT-R1  [^1] https://huggingface.co/datasets/IDEA-FinAI/Finance-R1-Reasoning, a synthetic dataset for stock prediction tasks.
Table 6: Extended performance metrics of financial large language models with additional models and datasets. Best results are marked in bold.
Task Dataset Metrics GPT-4o Yi-34B GLM-4-32B -0414 Mixtral 8x7B Qwen2.5 32B Qwen2.5 72B Xuanyuan 2.0 deepseek -v3 Touchstone GPT
Sentiment
Analysis
FPB Weighted-F1 0.8084 0.8652 0.8921 0.8510 0.8783 0.8815 0.8426 0.9014 0.8576
ACC 0.8093 0.8638 0.8905 0.8492 0.8761 0.8798 0.8403 0.8997 0.8557
Fiqa-SA Weighted-F1 0.8106 0.8321 0.8453 0.8257 0.8389 0.8421 0.8185 0.8512 0.8591
ACC 0.7702 0.8289 0.8402 0.8204 0.8337 0.8376 0.8123 0.8465 0.8638
Classification
Headlines Weighted-F1 0.7857 0.9521 0.9623 0.9387 0.9589 0.9610 0.9324 0.9684 0.9866
ACC 0.7931 0.9518 0.9619 0.9375 0.9582 0.9607 0.9312 0.9679 0.9866
FOMC Weighted-F1 0.6603 0.7824 0.8125 0.7653 0.8021 0.8089 0.7527 0.8237 0.8788
ACC 0.6794 0.7852 0.8183 0.7716 0.8079 0.8135 0.7591 0.8294 0.8790
lendingclub Weighted-F1 0.6730 0.8927 0.9215 0.8753 0.9128 0.9186 0.8624 0.9357 0.9783
MCC 0.1642 0.7825 0.8432 0.7521 0.8279 0.8375 0.7326 0.8674 0.9297
Entity
Extraction
NER Entity-F1 0.1800 0.4521 0.5832 0.4125 0.5527 0.5718 0.3924 0.6127 0.6993
Relation
Extraction
FinRE Relation-F1 0.1613 0.2854 0.3726 0.2543 0.3428 0.3615 0.2327 0.4128 0.5331
Multiple
Choice
CFA Weighted-F1 0.7700 0.6521 0.7125 0.6234 0.6928 0.7036 0.6057 0.7289 0.7497
ACC 0.7700 0.6528 0.7132 0.6241 0.6935 0.7043 0.6064 0.7296 0.7500
Summarization
EDTSUM Rouge-1 0.1675 0.3521 0.4128 0.3276 0.3987 0.4089 0.3124 0.4287 0.5254
Rouge-2 0.0556 0.2187 0.2875 0.2013 0.2739 0.2832 0.1924 0.3057 0.3446
Rouge-L 0.1069 0.2854 0.3521 0.2678 0.3389 0.3476 0.2543 0.3724 0.4705
BLEU 0.1192 0.3025 0.3687 0.2854 0.3562 0.3641 0.2728 0.3865 0.4512
Question
Answering
Finqa RMACC 0.1037 0.1682 0.1925 0.1543 0.1837 0.1891 0.1476 0.2014 0.2258
Convfinqa RMACC 0.2540 0.3827 0.4216 0.3652 0.4089 0.4173 0.3524 0.4385 0.5053
Stock Movement
Prediction
CMIN-CoT-R1 Weighted-F1 0.5025 0.4873 0.4982 0.4791 0.4924 0.4957 0.4738 0.5011 0.5136
ACC 0.5149 0.5027 0.5103 0.4975 0.5068 0.5092 0.4921 0.5138 0.5224
Appendix C Inference Template of Large Language Models
The Table.7 showcases how inference templates vary across different models. It is crucial to select the appropriate template for constructing correct inputs when inferring on the test sets of datasets. An incorrect template can significantly impair the performance of a model. We have observed that the underperformance of some large financial language models in some benchmarks is precisely due to not selecting the appropriate templates for evaluation. For more details of training and inference template, please refer to our open-source code repository on Github.